# Hidden memory and stochastic fluctuations in science

### v1

@@ Abstract
Understanding the statistical laws governing citation dynamics remains a fundamental challenge in network theory and the science of science. Citation networks typically exhibit in-degree distributions well approximated by log-normal distributions, yet they also display power-law behaviour in the high-citation regime, presenting an apparent contradiction that lacks a unified explanation. Here, we identify a previously unrecognised phenomenon: the variance of the logarithm of citation counts per unit time follows a power law with respect to time since publication, scaling as $t^{H}$. This discovery introduces a new challenge while simultaneously offering a crucial clue to resolving this discrepancy. We develop a stochastic model in which latent attention to publications evolves through a memory-driven process incorporating cumulative advantage. This process is characterised by the Hurst parameter $H$, derived from fractional Brownian motion, and volatility. Our framework reconciles this contradiction by demonstrating that anti-persistent fluctuations ($H<\tfrac{1}{2}$) give rise to log-normal citation distributions, whereas persistent dynamics ($H>\tfrac{1}{2}$) favour heavy-tailed power laws. Numerical simulations confirm our model’s explanatory and predictive power, interpolating between log-normal and power-law distributions while reproducing the $t^{H}$ law. Empirical analysis of arXiv e-prints further supports our theory, revealing an intrinsically anti-persistent nature with an upper bound of approximately $H=0.13$. By linking memory effects and stochastic fluctuations to broader network dynamics, our findings provide a unifying framework for understanding the evolution of collective attention in science and other attention-driven processes.
@@ I Introduction
The study of stochastic fluctuations in physics, originating from Einstein’s seminal investigation of Brownian motion [1], has reshaped our understanding of the natural world. Often unpredictable at the microscopic level, these dynamics have revealed fundamental parameters and governing principles through collective behaviour. This realisation has led to the recognition that many physical phenomena can be described within a universal framework, independent of their microscopic details. These insights have been pivotal in advancing non-equilibrium statistical mechanics—encompassing nonlinear dynamics and network theory—and have influenced chaos and fractal theory [2, 3]. Beyond physics and mathematics, they have helped explain biological morphogenesis [4, *Keller71], the structural and dynamic evolution of large-scale networks such as the Internet, and complex social phenomena [6, *Albert02].
Among the large-scale networks that exemplify these principles, a particularly relevant example arises in academia. In this domain, the scholarly practice of citation, where researchers systematically reference prior work to position their own findings, gives rise to a citation network [8]. This network is constructed through the cumulative assembly of references and bibliographies across a vast collection of publications. Over the course of scientific history, this citation network has evolved into a substantial directed graph with a distinctive structure.
@@
Another fundamental aspect of citation network research concerns its temporal evolution, specifically how the number of citations to individual publications changes over time [8, 27, *Nakamoto88, *Lariviere08, *Parolo15, *Yin17, 32, 33, 34, 22]. Most publications receive few or no citations [8, 35, 10]. Among those that do, some gain recognition only long after publication [17, 36, *Ke15], while others continue accumulating citations indefinitely. Each case follows a distinct citation pattern, yet collectively they exhibit a characteristic ‘jump-and-decay’ trend: citations rise sharply after publication, peak briefly, and then gradually decline, asymptotically approaching zero with a slower descent than the initial rise.
@@
A fundamental mechanism underlying citation dynamics, widely recognised as a key empirical observation, is cumulative advantage [9]. This principle states that entities with an initial advantage tend to accumulate even more over time. In networks, preferential attachment [38, *Barabasi02, *Jeong03, *Newman09] formalises this idea by favouring nodes with a greater number of links. When applied linearly, it leads to a power-law distribution. Furthermore, Gibrat’s law [42] extends linear preferential attachment to a continuous, stochastic setting. Historically, these concepts have been framed as the Yule process [43, *Simon55] or, more commonly, the Matthew effect, often summarised as the ‘rich get richer’ phenomenon.
A key question in citation network research is whether a theoretical framework can be developed that accommodates diverse temporal citation trajectories at the individual level while collectively reproducing the characteristic in-degree distribution and incorporating preferential attachment. Addressing this question requires recognising a previously overlooked empirical observation: the standard deviation (SD) of the logarithm of citation counts evolves as a power law with respect to time since publication (the $t^{H}$ law; see Section II). Existing theories of citation dynamics do not naturally account for this phenomenon.
To bridge this gap, we propose a theoretical model that integrates and explains various empirical findings within a unified mathematical framework (Section III). The validity of our theory is first assessed through numerical simulations, evaluating its explanatory power and predictive capability (Section IV). It is then tested against empirical data to determine its effectiveness in describing real-world citation networks (Section V). Finally, we examine the broader implications of our framework (Section VI). This approach provides new insights into why citation network in-degree distributions are often well approximated by log-normal distributions while simultaneously exhibiting power-law behaviour in high-citation regimes. It also reveals an intrinsic fractal structure associated with citation trajectories. At the core of our theory lies the fundamental principle introduced at the outset: the science of stochastic fluctuations.
@@ II.1 Scientometric Engineering
A similar principle governs collective attention to science, as reflected in citation patterns, where both a predictable, controllable component and an inherently random component coexist. In this context, the ‘die’ represents not an individual publication but a broader entity, such as science as a whole or a specific research community in disciplines like physics or mathematics, where potential attention is distributed. The deterministic aspect of citation dynamics is shaped by quantifiable factors, including whether the author is an established scholar, the prominence of the publication venue (e.g., journal impact factor), prior citations and attention received, and intrinsic attributes such as fitness or novelty, provided they are properly quantified [34, 45, *Stegehuis15, *Cao16]. In contrast, stochastic elements arise from interactions with the citing community—researchers who may reference the work. These factors include cognitive limitations, resource constraints, time pressures, and the serendipitous discovery of new connections to other studies post-publication. Such unpredictable influences contribute to the inherent stochasticity of the citation process.
@@
More recently, a comparable strategy has been proposed for elucidating citation dynamics [23]. This work hypothesised that citation dynamics could similarly be modelled using an SDE, incorporating both deterministic and stochastic components, as in the Black–Scholes model. It led to the formulation of an SDE for citation dynamics based on a generalised version of Black–Scholes’ geometric Brownian motion. Yet, a crucial empirical phenomenon in citation dynamics, while hinted at in that study [23, Appendix C, Suppl. Fig. S9], was not explicitly emphasised [53]. As detailed below, this characteristic concerns the relationship between citation counts per unit time and the elapsed time since publication.
@@ II.2 The $\bm{t^{H}}$ law
Figure 1 illustrates the relationship between the SD of the logarithm of annual citation counts ($c$) and the number of years since posting ($t$) for e-prints published on arXiv (https://arxiv.org) in 2001 ($N=39{,}106$). As is evident from the figure, SD increases with $t$ following a power-law relationship with a positive exponent $H$:
@@
This relationship holds consistently across publications from different publication years, as well as when the analysis is confined to specific research fields rather than the entire arXiv. It also remains robust when alternative methods of conceptualising elapsed time are employed [53]. Given this apparent universality, this paper refers to Eq. (1) as the $t^{H}$ law. For the dataset in Fig. 1, a linear regression analysis estimates the exponent as $\hat{H}\approx 0.13\pm 0.016$ $(p<0.001)$. A more detailed analysis and discussion are presented later in Section V; for now, the empirical findings above already provide strong motivation for developing a theoretical framework with enhanced explanatory power.
Figure 1: The relationship between the standard deviation (SD) of the logarithm of annual citation counts ($c$) and the number of years since publication ($t$). The data are based on [54], focusing on e-prints published on arXiv in 2001 ($N=39{,}106$).
A previous study [23], using the same dataset as Fig. 1, concluded that the cumulative citation distribution for arXiv e-prints is better characterised by a log-normal distribution than a (shifted) power law. Notably, the shape parameter ($\sigma$) of the log-normal distribution does not directly correspond to $\operatorname{SD}[\ln c]$ in Eq. (1). This discrepancy arises because $\sigma$ characterises the distribution of cumulative citation counts over a publication’s lifetime or a long-term period of interest, whereas $\operatorname{SD}[\ln c]$ reflects citation counts within a short, fixed time window (in this case, each year). The two quantities are connected through a time-integrated relationship, both increasing monotonically with time since publication. Indeed, an empirical relationship, $\sigma\sim t^{H^{\prime}}$ (where $H^{\prime}=\text{const.}$.), has been observed [23]. This trend can be interpreted as an effective law arising from the ‘convolution’ of the $t^{H}$ law in Eq. (1).
In previous studies, when the cumulative distribution was identified as log-normal, the parameter $\sigma$ was typically treated as a static quantity. Although its estimated value varies across datasets and research fields, it is generally reported to range from 1 to 1.4 [18, *Evans12, *Chatterjee16, *DAngelo19, 32, 22, 23]. However, these studies did not investigate the temporal behaviour of $\sigma$, especially its consistent increase with time since publication. Consequently, understanding the mechanisms underlying the $t^{H}$ law and developing a theoretical model that naturally reproduces this relationship remain important open problems [53], which we address below.
@@ III.1 Heuristic considerations
Second, let us revisit the left-hand side of the $t^{H}$ law (1) and refine the definition of the quantity $c$ with higher temporal resolution. To this end, let $C(t)$ denote the cumulative citation count at time $t$, $\Delta t$ a sufficiently small time unit (one year in Fig. 1), and $\Delta C(t)=C(t+\Delta t)-C(t)$ the change over this interval. The citation rate is then given by $c(t;\Delta t)=\Delta C(t)/\Delta t$. It is important to note that $C(t)$ does not represent the citation count of a single publication at time $t$ but rather the aggregate of discrete citation events reflecting attention to the collective body of publications. Consequently, $C(t)$ is a monotonically increasing function that takes non-negative integer values, forming a piecewise constant curve with numerous localised steps. If the discrete jumps at each time $t=t_{k}$ are denoted by $a_{k}\in\mathbb{N}$ and $\Theta(x)$ represents the Heaviside step function, then $C(t)$ can be expressed as $C(t)=\sum_{k}a_{k}\Theta(t-t_{k})$.
Now, consider the limit $\Delta t\to 0$. At the discrete jump points in the graph of $C(t)$, the function is not differentiable in the usual sense; instead, its derivative behaves analogously to Dirac’s $\delta$-function. Specifically, the instantaneous attention at time $t$ can be expressed as:
@@
If we refine the temporal resolution of attention history curves—typically observed as jump-and-decay patterns on an annual scale—to the level of instantaneous values and aggregate them across all relevant publications, the resulting function, though continuous, appears highly non-smooth. This function represents a superposition of numerous sharp spikes, forming a randomly jagged trajectory reminiscent of Brownian motion. In essence, the trajectory undergoes random fluctuations, while its overall trend follows a jump-and-decay pattern.
III.2 The Ansatze
Building on the above mathematical and physical insights, the core ansatze for modelling citation dynamics are as follows:
@@ III.2 The Ansatze
    Here, $\mathcal{A}(t)$ represents the average or expected trajectory of $X(t)$, while the stochastic component $\mathcal{B}(t)$ captures random fluctuations.
@@
Several remarks follow regarding these ansatze. First, with respect to Ansatz (i), the relationship between the latent attention function $X(t)$ and its integral over the interval $[0,T]$, namely $C(T)$, will be discussed in detail later. In particular, we will clarify how the continuous attention variable relates to the discrete citation count (see Eq. (7) and the discussion in Section VI.3).
@@
For Ansatz (iii), we note a few points about the connection between Gibrat’s law and preferential attachment. While both involve multiplicative growth processes, there are some key differences. Preferential attachment typically describes discrete nodes in a network acquiring new links, whereas Gibrat’s law is usually applied to continuous variables, such as wealth or firm sizes—and here, to latent attention $X(t)$. Moreover, preferential attachment, though probabilistic at the microscopic level, effectively follows a global deterministic rule that yields a power law distribution if the attachment function is linear [38, *Barabasi02, *Jeong03, *Newman09]. By contrast, Gibrat’s law under stochastic fluctuations in the proportionality factor generates a log-normal distribution via geometric Brownian motion [42, 23]. This feature will be crucial when we discuss the distribution of $C(T)$ in later sections.
Lastly, it is worth noting that the above set of ansatze applies not only to citations but also to other forms of attention process. In this context, citation can be replaced with similar indices that reflect the collective attention directed towards a given entity within a community. See also the discussion in Section VI.1 [53].
@@ III.3 Key Properties of Fractional Brownian Motion
Fractional Brownian motion, denoted as $\{B_{H}(t)\}_{t\geq 0}$ with the Hurst parameter $H\in(0,1)$, is a class of non-Markovian and non-martingale, continuous-time Gaussian stochastic processes with stationary increments. It satisfies
@@
From Eqs. (4, 5), it can be shown that for the fractional Brownian motion with $H\neq\tfrac{1}{2}$, a process after a given time $t$ depends not only on the situation at time $t$ but also on the entire past history of the process up to time $t$. Specifically, the fractional Brownian motion increment $\Delta B_{H}(t,s)\coloneqq B_{H}(t)-B_{H}(s)$ with $0\leq s\leq t$ has the moment properties that $\mathbb{E}[\Delta B_{H}(t,s)]=0$ and $\mathbb{E}[\Delta B_{H}(t,s)^{2}]=\left|t-s\right|^{2H}$, and therefore, the covariance of two non-overlapping increments is given by
@@
Consequently, if $H\in(\tfrac{1}{2},1)$, the increments of fractional Brownian motion exhibit positive correlation (persistence), meaning a tendency for future increments to continue in the same direction as past increments, and the resulting motion is superdiffusive. Conversely, if $H\in(0,\tfrac{1}{2})$, the increments are negatively correlated (anti-persistence), meaning a tendency for future increments to reverse the direction of past increments, producing subdiffusive motion. Roughly speaking, the anti-persistent case corresponds to more chaotic behaviour, while the persistent case reflects more structured and disciplined behaviour.
@@
$D=2-H\,.$ (6)
As can be seen from this formula, the closer the value of $H$ is to 1—that is, the stronger the persistent tendency—the trajectory approaches a smoother line. Conversely, the closer $H$ is to 0—that is, the stronger the anti-persistent tendency—the trajectory becomes more jagged and complex in shape.
@@
The distinctive and versatile nature of fractional Brownian motion makes it a compelling model for phenomena exhibiting both short-range and long-range dependence across various fields, including biology [57, *Jeon11], hydrology [59], telecommunications and network traffic [60, *Abry00, *Mikosch02], and economics and finance [63, *Comte98, 65]. Indeed, numerous studies have demonstrated that many phenomena and challenges in these domains are better captured by fractional Brownian motion, which incorporates memory effects over time.

### v2

# Fractional stochastic model of citation dynamics with memory and volatility

@@ Abstract
Understanding the statistical laws governing citation dynamics remains a fundamental challenge in network theory and the science of science. Citation networks typically exhibit in-degree distributions well approximated by log-normal distributions yet also display power-law behaviour in the high-citation regime—an apparent contradiction lacking a unified explanation. Here we identify a previously unrecognised phenomenon: the variance of the logarithm of citation counts per unit time follows a power law with respect to time ($t$) since publication, scaling as $t^{H}$, with $H$ constant. This discovery introduces a new challenge while simultaneously offering a crucial clue to resolving this discrepancy. We develop a stochastic model in which latent attention to publications evolves through a memory-driven process with cumulative advantage, modelled as fractional Brownian motion with Hurst parameter $H$ and volatility. We show that antipersistent fluctuations in attention ($H<1/2$) yield log-normal citation distributions, whereas persistent attention dynamics ($H>1/2$) favour heavy-tailed power laws, thus resolving the log-normal–power-law contradiction. Numerical simulations confirm both the $t^{H}$ law and the transition between regimes. Empirical analysis of arXiv e-prints indicates that the latent attention process is intrinsically antipersistent ($H\approx 0.13$). By linking memory effects and stochastic fluctuations in attention to broader network dynamics, our findings provide a unifying framework for understanding the evolution of collective attention in science and other attention-driven processes.
DOI: 10.1103/l2xd-43n9
@@ I Introduction
The study of stochastic fluctuations in physics, originating from Einstein’s seminal investigation of Brownian motion [1], has reshaped our understanding of the natural world. Often unpredictable at the microscopic level, these dynamics have revealed fundamental parameters and governing principles through collective behaviour. This realisation has led to the recognition that many physical phenomena can be described within a universal framework, independent of their microscopic details. These insights have been pivotal in advancing nonequilibrium statistical mechanics—encompassing nonlinear dynamics and network theory—and have influenced chaos and fractal theory [2, 3]. Beyond physics and mathematics, they have helped explain biological morphogenesis [4, *Keller71], the structural and dynamic evolution of large-scale networks such as the Internet, and complex social phenomena [6, *Albert02].
Among the large-scale networks that exemplify these principles, a particularly relevant example arises in academia itself. In this domain, the scholarly practice of citation, where researchers systematically reference prior work to position their own findings, gives rise to a citation network [8]. This network is constructed through the cumulative assembly of references and bibliographies across a vast collection of publications. Over the course of scientific history, this citation network has evolved into a substantial directed graph with a distinctive structure.
@@
Another fundamental aspect of citation network research concerns its temporal evolution, specifically how the number of citations to individual publications changes over time [8, 27, *Nakamoto88, *Lariviere08, *Parolo15, *Yin17, 32, 33, 34, 22]. Most publications receive few or no citations [8, 35, 10]. Among those that do, some gain recognition only long after publication [17, 36, *Ke15], while others continue accumulating citations indefinitely. Each case follows a distinct citation pattern yet collectively they exhibit a characteristic ‘jump-and-decay’ trend: citations rise sharply after publication, peak briefly, and then gradually decline, asymptotically approaching zero with a slower descent than the initial rise.
@@
A paradigmatic mechanism often invoked to explain citation dynamics is cumulative advantage [9], which has played a foundational role in modelling efforts, despite its known limitations in reproducing detailed structures of real networks. This principle refers to the idea that entities with an initial advantage—such as early citations—tend to attract further attention, thereby accumulating disproportionately more citations over time. In networks, preferential attachment [38, *Barabasi02, *Jeong03, *Newman09] formalises this idea by favouring nodes with a greater number of links. When applied linearly, it leads to a power-law distribution. Furthermore, Gibrat’s law [42] extends linear preferential attachment to a continuous, stochastic setting. Historically, these concepts have been framed as the Yule process [43, *Simon55] or, more commonly, the Matthew effect, often summarised as the ‘rich get richer’ phenomenon.
A key question in citation network research is whether a theoretical framework can be developed that accommodates diverse temporal citation trajectories at the individual level while collectively reproducing the characteristic in-degree distribution and incorporating preferential attachment. Addressing this question requires recognising a previously overlooked empirical observation: the standard deviation (SD) of the logarithm of citation counts evolves as a power law with respect to time since publication (the $t^{H}$ law; see Sec. II). Existing theories of citation dynamics do not naturally account for this phenomenon.
To bridge this gap, we propose a theoretical model that unifies and explains a range of empirical findings within a coherent mathematical framework (Sec. III). The validity of our theory is first examined through numerical simulations, which assess both its explanatory strength and predictive capability (Sec. IV). We then test the model against empirical data to evaluate its effectiveness in describing real-world citation networks (Sec. V). Finally, we explore the broader implications of our framework (Sec. VI).
Our approach offers fresh insight into why in-degree distributions in citation networks are often well approximated by log-normal forms while simultaneously exhibiting power-law behaviour in the high-citation regime. It also uncovers an intrinsic fractal structure underlying citation trajectories. At the heart of our theory lies a fundamental principle introduced at the outset: the science of stochastic fluctuations.
@@ II.1 Scientometric Engineering
A similar principle governs collective attention to science, as reflected in citation patterns, where both a predictable, controllable component and an inherently random component coexist. In this context, the ‘die’ represents not an individual publication but a broader entity, such as science as a whole or a specific research community in disciplines like physics or mathematics, where potential attention is distributed. The deterministic aspect of citation dynamics is shaped by quantifiable factors, including whether the author is an established scholar, the prominence of the publication venue (e.g., journal impact factor), prior citations and attention received, and intrinsic attributes such as fitness or novelty, provided they are properly quantified [34, 45, *Stegehuis15, *Cao16]. In contrast, stochastic elements arise from interactions with the citing community—researchers who may reference the work. These factors include cognitive limitations, resource constraints, time pressures, and the serendipitous discovery of new connections to other studies postpublication. Such unpredictable influences contribute to the inherent stochasticity of the citation process.
@@
More recently, a comparable strategy has been proposed for elucidating citation dynamics [23]. This work hypothesised that citation dynamics could similarly be modelled using an SDE, incorporating both deterministic and stochastic components, as in the Black–Scholes model. It led to the formulation of an SDE for citation dynamics based on a generalised version of Black–Scholes’ geometric Brownian motion. Yet a crucial empirical phenomenon in citation dynamics, while hinted at in that study (Appendix C, Supplemental Fig. S9 in Ref. [23]), was not explicitly emphasised [53]. As detailed below, this characteristic concerns the relationship between citation counts per unit time and the elapsed time since publication.
@@ II.2 The $\bm{t^{H}}$ law
Figure 1 illustrates the relationship between the SD of the logarithm of annual citation counts ($c$) and the number of years since posting ($t$) for e-prints published on arXiv [54] in 2001 (N=39,106N=39{,}106). As is evident from the figure, SD increases with $t$ following a power-law relationship with a positive exponent $H$:
@@
This relationship holds consistently across publications from different publication years, as well as when the analysis is confined to specific research fields rather than the entire arXiv. It also remains robust when alternative methods of conceptualising elapsed time are employed [53]. Given this apparent universality, this paper refers to Eq. (1) as the $t^{H}$ law. For the dataset in Fig. 1, a linear regression analysis estimates the exponent as $\hat{H}\approx 0.13\pm 0.016$ $(p<0.001)$. A more detailed analysis and discussion are presented later in Sec. V; for now, the empirical findings above already provide strong motivation for developing a theoretical framework with enhanced explanatory power.
Figure 1: The relationship between the standard deviation (SD) of the logarithm of annual citation counts ($c$) and the number of years since publication ($t$). The data are based on Ref. [55], focusing on e-prints published on arXiv in 2001 (N=39,106N=39{,}106).
A previous study [23], using the same dataset as Fig. 1, concluded that the cumulative citation distribution for arXiv e-prints is better characterised by a log-normal distribution than a (shifted) power law. Notably, the shape parameter ($\sigma$) of the log-normal distribution does not directly correspond to $\operatorname{SD}[\ln c]$ in Eq. (1). This discrepancy arises because $\sigma$ characterises the distribution of cumulative citation counts over a publication’s lifetime or a long-term period of interest, whereas $\operatorname{SD}[\ln c]$ reflects citation counts within a short, fixed time window (in this case, each year). The two quantities are connected through a time-integrated relationship, both increasing monotonically with time since publication. Indeed, an empirical relationship, $\sigma\sim t^{H^{\prime}}$, with $H^{\prime}$ constant, has been observed [23]. This trend can be interpreted as an effective law arising from the ‘convolution’ of the $t^{H}$ law in Eq. (1).
In previous studies, when the cumulative distribution was identified as log-normal, the parameter $\sigma$ was typically treated as a static quantity. Although its estimated value varies across datasets and research fields, it is generally reported to range from 1 to 1.4 [18, *Evans12, *Chatterjee16, *DAngelo19, 32, 22, 23]. However, these studies did not investigate the temporal behaviour of $\sigma$, especially its consistent increase with time since publication. Consequently, understanding the mechanisms underlying the $t^{H}$ law and developing a theoretical model that naturally reproduces this relationship remain important open problems, which we address below.
@@ III.1 Heuristic considerations
Second, let us revisit the left-hand side of the $t^{H}$ law (1) and refine the definition of the quantity $c$ with higher temporal resolution. To this end, let $C(t)$ denote the cumulative citation count at time $t$, $\Delta t$ a sufficiently small time unit (one year in Fig. 1), and $\Delta C(t)=C(t+\Delta t)-C(t)$ the change over this interval. The citation rate is then given by $c(t;\Delta t)=\Delta C(t)/\Delta t$. It is important to note that $C(t)$ does not represent the citation count of a single publication at time $t$ but rather the aggregate of discrete citation events reflecting attention to the collective body of publications. Consequently, $C(t)$ is a monotonically increasing function that takes nonnegative integer values, forming a piecewise constant curve with numerous localised steps. If the discrete jumps at each time $t=t_{k}$ are denoted by $a_{k}\in\mathbb{N}$ and $\Theta(x)$ represents the Heaviside step function, then $C(t)$ can be expressed as $C(t)=\sum_{k}a_{k}\Theta(t-t_{k})$.
Now consider the limit $\Delta t\to 0$. At the discrete jump points in the graph of $C(t)$, the function is not differentiable in the usual sense; instead, its derivative behaves analogously to Dirac’s $\delta$ function. Specifically, the instantaneous attention at time $t$ can be expressed as:
@@
If we refine the temporal resolution of attention history curves—typically observed as jump-and-decay patterns on an annual scale—to the level of instantaneous values and aggregate them across all relevant publications, then the resulting function, though continuous, appears highly nonsmooth. This function represents a superposition of numerous sharp spikes, forming a randomly jagged trajectory reminiscent of Brownian motion. In essence, the trajectory undergoes random fluctuations, while its overall trend follows a jump-and-decay pattern.
III.2 The Ansätze
Building on the above mathematical and physical insights, the core Ansätze for modelling citation dynamics are as follows:
@@ III.2 The Ansätze
    Here $\mathcal{A}(t)$ represents the average or expected trajectory of $X(t)$, while the stochastic component $\mathcal{B}(t)$ captures random fluctuations.
@@
Several remarks follow regarding these Ansätze. First, with respect to Ansatz (i), the relationship between the latent attention function $X(t)$ and its integral over the interval $[0,T]$, namely $C(T)$, will be discussed in detail later. In particular, we will clarify how the continuous attention variable relates to the discrete citation count (see Eq. (7) and the discussion in Sec. VI.3).
@@
For Ansatz (iii), we note a few points about the connection between Gibrat’s law and preferential attachment. While both involve multiplicative growth processes, there are some key differences. Preferential attachment typically describes discrete nodes in a network acquiring new links, whereas Gibrat’s law is usually applied to continuous variables, such as wealth or firm sizes—and here to latent attention $X(t)$. Moreover, preferential attachment, though probabilistic at the microscopic level, effectively follows a global deterministic rule that yields a power law distribution if the attachment function is linear [38, *Barabasi02, *Jeong03, *Newman09]. By contrast, Gibrat’s law under stochastic fluctuations in the proportionality factor generates a log-normal distribution via geometric Brownian motion [42, 23]. This feature will be crucial when we discuss the distribution of $C(T)$ in later sections.
Last, it is worth noting that the above set of Ansätze applies not only to citations but also to other forms of attention process. In this context, citation can be replaced with similar indices that reflect the collective attention directed towards a given entity within a community. See also the discussion in Sec. VI.1 [53].
@@ III.3 Key properties of fractional Brownian motion
Fractional Brownian motion, denoted as $\{B_{H}(t)\}_{t\geq 0}$ with the Hurst parameter $H\in(0,1)$, is a class of non-Markovian and nonmartingale, continuous-time Gaussian stochastic processes with stationary increments. It satisfies
@@
From Eqs. (4) and (5), it can be shown that for the fractional Brownian motion with $H\neq\tfrac{1}{2}$, a process after a given time $t$ depends not only on the situation at time $t$ but also on the entire past history of the process up to time $t$. Specifically, the fractional Brownian motion increment $\Delta B_{H}(t,s)\coloneqq B_{H}(t)-B_{H}(s)$ with $0\leq s\leq t$ has the moment properties that $\mathbb{E}[\Delta B_{H}(t,s)]=0$ and $\mathbb{E}[\Delta B_{H}(t,s)^{2}]=\left|t-s\right|^{2H}$, and therefore, the covariance of two nonoverlapping increments is given by
@@
Consequently, if $H\in(\tfrac{1}{2},1)$, then the increments of fractional Brownian motion exhibit positive correlation (persistence), meaning a tendency for future increments to continue in the same direction as past increments, and the resulting motion is superdiffusive. Conversely, if $H\in(0,\tfrac{1}{2})$, then the increments are negatively correlated (antipersistence), meaning a tendency for future increments to reverse the direction of past increments, producing subdiffusive motion. Roughly speaking, the antipersistent case corresponds to more chaotic behaviour, while the persistent case reflects more structured and disciplined behaviour.
@@
D=2H.D=2-H\,. (6)
As can be seen from this formula, the closer the value of $H$ is to 1—that is, the stronger the persistent tendency—the trajectory approaches a smoother line. Conversely, the closer $H$ is to 0—that is, the stronger the antipersistent tendency—the trajectory becomes more jagged and complex in shape.
@@
The distinctive and versatile nature of fractional Brownian motion makes it a compelling model for phenomena exhibiting both short-range and long-range dependence across various fields, including biology [58, *Jeon11], hydrology [60], telecommunications and network traffic [61, *Abry00, *Mikosch02], and economics and finance [64, *Comte98, 66]. Indeed, numerous studies have demonstrated that many phenomena and challenges in these domains are better captured by fractional Brownian motion, which incorporates memory effects over time.